# Image-Based Roadmaps for Vision-Only Planning and Control of Robotic Manipulators

### v1

@@ Abstract
This work presents a motion planning framework for robotic manipulators that computes collision-free paths directly in image space. The generated paths can then be tracked using vision-based control, eliminating the need for an explicit robot model or proprioceptive sensing. At the core of our approach is the construction of a roadmap entirely in image space. To achieve this, we explicitly define sampling, nearest-neighbor selection, and collision checking based on visual features rather than geometric models. We first collect a set of image-space samples by moving the robot within its workspace, capturing keypoints along its body at different configurations. These samples serve as nodes in the roadmap, which we construct using either learned or predefined distance metrics. At runtime, the roadmap generates collision-free paths directly in image space, removing the need for a robot model or joint encoders. We validate our approach through an experimental study in which a robotic arm follows planned paths using an adaptive vision-based control scheme to avoid obstacles. The results show that paths generated with the learned-distance roadmap achieved 100% success in control convergence, whereas the predefined image-space distance roadmap enabled faster transient responses but had a lower success rate in convergence.
@@ I Introduction
The goal of our research is to push the boundaries of purely vision-based control and motion planning for robotic manipulators, by decreasing reliance on explicit robot modeling or proprioceptive sensing. In doing so, we strive to use natural visual features along the robot’s body in image space, without attaching any external markers. While the works in [15, 16, 17] provided algorithms to track natural keypoints and use them to achieve vision-based model-free control with decent transient responses, these algorithms are only designed to run in obstacle-free space and do not provide motion planning capability to avoid obstacles in the scene.
@@
In this work, we tackle the problem of collision free path planning for purely vision-based robot control: we develop a planning and control scheme that does not rely on a robot model or on-board sensors (e.g. joint encoders) in runtime. As such, this strategy is especially useful for (but not limited to) soft/underactuated/inexpensive robots that are hard to model and may not have reliable proprioceptive sensing.
@@
This dataset was used to develop a simple neural network that approximates joint displacements based on keypoint locations in image space. These distance metrics were integrated into the roadmap construction. Once the roadmaps are constructed, polygon-based collision checking and A* search are employed to ensure collision-free paths. These two approaches have different implications for vision-based robot control. In a nutshell, we observed that utilizing estimated joint distances in the roadmap results in smoother and more accurate tracking of the generated path, allowing the robot to stay in its defined workspace and avoid obstacles. In contrast, roadmaps based on Euclidean distances in image space can offer faster transient responses, albeit with potentially less reliable tracking. Our experiments demonstrate these aspects both in the presence and absence of obstacles.
@@ II Related Work
Among related works, the approaches proposed in [28] and [29] are the most closely aligned with ours. In [28], the authors use sampling-based planning directly from images by leveraging learned forward propagation models, custom distance metrics, and collision checkers. While effective, their method requires extensive simulation-based training, which introduces a significant sim-to-real gap. In contrast, our method uses a limited dataset collected entirely from real robots and constructs a configuration-space roadmap without relying on simulation. Similarly, [29] combines image features with a robot model to train a neural network for planning a path. Unlike their approach, our method eliminates the need for a robot model, relying solely on image features for motion planning.
@@ III-A Model-based Motion Planning
$\mathcal{X}_{\text{obs}}=\{x\in\mathcal{X}|R(x)\bigcap\mathcal{W}\neq\emptyset\}$ (2)
where $R(x)\subseteq\mathbb{R}^{3}$ is the set of all the points of the robot in the workspace at state x [^1] This equation only encodes collisions but joint/kinematic constraints can be encoded in a similar manner. Then let, $\mathcal{X}_{\text{free}}=\mathcal{X}\setminus\mathcal{X}_{\text{obs}}$ denote the collision free state-space. Also, let $x_{\text{start}}\in\mathcal{X}_{\text{free}}$ and $X_{\text{goal}}\subseteq\mathcal{X}_{\text{free}}$ denote the start state and goal regions respectively.
Motion Planning Problem: Given the motion planning tuple $(\mathcal{X},R,\mathcal{W},\mathcal{U},f,x_{\text{start}},\mathcal{X}_{G})$ find a time $\mathcal{T}$ and a set of controls $u:[0,\mathcal{T}]\rightarrow\mathcal{U}$ such that the motion described by Eq. 1 satisfies $x(0)=x_{\text{start}}$, $x(\mathcal{T})\in\mathcal{X}_{G}$ and $x(t)\in\mathcal{X}_{\text{free}}$.
III-B Vision-Only Motion Planning
@@
$I_{\text{obs}}=\{\mathbb{K}\in I|RI(\mathbb{K})\bigcap\mathcal{IO}\neq\emptyset\}$ (3)
@@
Here we note that even if the underlying function f is fully-integrable (holonomic-system) if $dim(I)>dim(\mathcal{X})$ the equivalent system dynamics equations $g$ will be non-holonomic as there will be paths in the higher dimensional image-space $I$ that cannot be followed by the robot. In our approach we consider this issue, and propose a way to produce paths that approximately satisfy these constraints. In the first image of Fig. 2, the grey region illustrates the discretized representation of $I$, as each configuration is captured as an image frame, resulting in a finite set of $\mathbb{K}$ sampled at a specific frame rate. The second image of Fig. 2 highlights similar discretized representation of $I_{\text{free}}$.
@@ IV Methodology
To compute the path, which is the main contribution of this work, we propose to use probabilistic road-map planner (prm) by adapting its subroutines to operate directly in $I$-space. Specifically, we opted to adapt the lazy-prm [32] planner due to its compatibility with our particular requirements. However, any sampling-based planner which relies on the the same subroutines could be used.
@@
Similar to prm, lazy-prm operates in two distinct phases, the building-phase Alg. 1 and the query-phase Alg. 2. In the building-phase, a roadmap is built without performing any collision checking. In the query-phase, the roadmap is utilized to find a potential path for a given start and goal. The key difference between lazy-prm and prm lies in their approach to collision checking. While prm verifies the collision status of all edges in the roadmap during the building phase, lazy-prm reverses the order. It first performs the search query and only checks collisions for the found path. If any edges are found to be in collision, the roadmap is updated, and another search query is executed until a valid path is discovered.
We selected probabilistic roadmap method for its efficiency in multi-query scenarios, where a single precomputed graph can handle multiple start and goal configurations, and used the lazy version, as we only have very few varying obstacles between queries. Additionally, its ability to operate in high-dimensional spaces makes it particularly well-suited for handling non-traditional representations of configuration space, such as the visual keypoints in $I$-space used in our approach. We first describe how lazy-prm works and then we describe how we modified the necessary subroutines to make it work in $I$-space.
The building-phase (Alg. 1) works with the following procedure. First, in line 4, a sample is generated and added in graph $G$. Then the k nearest nearest neighbors are found by using a distance defined in $I$ (line 7) and are connected with edges. This continues until $N$ nodes are in the graph $G$.
@@
During the query-phase (Alg. 2 a new motion planning problem is solved. Given a $\mathbb{K}_{\text{start}}$ and $\mathbb{K}_{\text{goal}}$ they are added in the graph, and connected with their nearest-neighbors. Then a graph search algorithm e.g., A* is used to find a path. If edges of the path are in-collision they are updated accordingly in the roadmap, and the process repeats until a collision free-path is found. The three operations sample, k-nearest, coll_free, for Alg. 1 in lines 4, 7 and Alg. 2 typically require a model for the robot.
@@
sample: This function typically samples the configuration space uniformly. However, in the vision-only setting, we can’t directly sample in $I$-space as we don’t have the model of the robot. Since $dim(I)>dim(\mathcal{X})$ the keypoint vectors (image state $\mathbb{K}$) that correspond to actual configurations of the robot, will lie in a lower-dimensional (equal to $dim(\mathcal{X})$) manifold in $I$-space. Thus randomly sampling $I$-space will have 00 probability of sampling a valid configuration that lies on the valid manifold [22]. We describe how to address this issues in subsection IV-A, by collecting and storing valid samples directly from the real robot.
k-nearest: This function usually relies on a distance defined in $\mathcal{X}$ and finds the nearest configurations that can be connected. However, since our representation in $I$-space is now a non-holonomic system, defining this distance is very challenging, as a straight line path in $I$-space defined by a simple Euclidean distance, might not be accurately followable by a controller. To mitigate this, we describe a learning-based approach to estimate the unknown joint-distance, in subsection IV-B.
coll_free: This function checks if there is a collision for an edge in $\mathcal{X}$. Again this typically requires the model $R(x)$ for the robot. In the vision only case, we propose simple yet successful method to do collision checking with an $RI(x)$ directly in $I$-space subsection IV-C.
In the next section we describe our proposed method, for each, of the aforementioned subroutines. Fig. 3 visually describes all the steps of the visual motion planning framework.
@@
$v_{j}=\min\left(\frac{M_{j}}{res\cdot t},v_{\text{max}}\right),$ (5)
where, $M_{j}$ is the motion range or difference of limits of joint $j$, $res$, is the number of discrete steps used to traverse $M_{j}$, $t$ is the duration allocated to complete each step, and $v_{\text{max}}$ is the maximum allowable velocity for joint $j$. By dividing the motion range of each joint into uniform increments based on $res$, we ensure that the robot systematically explores all possible configurations within its workspace. The resulting dataset of image states $\mathbb{K}$ is thus evenly distributed across the image space $I$, enabling robust coverage and accurate representation of the robot’s motion capabilities.
@@
$\mathbb{K}_{n}=K\cdot T_{cw}\cdot x_{n}$ (6)
where $K$ is the camera intrinsic matrix, $T_{cw}$ is the camera extrinsics matrix derived from calibration processes described in [33, 34], and $x_{n}$ is the 3D configuration robot in workspace. This transformation projects the 3D configuration $x_{n}$, into their corresponding 2D (pixel) projection in the image described in [35], resulting in a set of $\mathds{k}$ for each image state $\mathbb{K}_{n}$ The process captures the robot’s motion across its visible workspace and creates a comprehensive dataset of $\mathbb{K}$, representing close to all feasible configurations in image space. Dataset of $\mathbb{K}$ can also be collected by following the process describe in [17].
The Keypoint Dataset Samples section of Fig. 3, represents this part of the workflow. The grey area in the left image of Fig. 2 illustrates how $\mathbb{K}$ are distributed within image space with each frame consisting of a set of $\mathds{k}$ or keypoints. This collection process yields a large dataset of $\mathbb{K}$ that can be used to enable efficient roadmap construction for vision-based motion planning.
IV-B k-nearest: Learned and Image-Space Metrics
@@
  • Learned distance, where the distance is learned by a neural network, trained to predict joint displacements between two image states. The input to the network is a pair of image states $\mathbb{K}_{1}$ and $\mathbb{K}_{2}$ and and the output is an estimated joint displacement required to transition between them:
    $\text{dist}_{learned}()\leftarrow\text{NN}(\mathbb{K}_{1},\mathbb{K}_{2})$ (7)
@@
    $\text{dist}_{image}()\leftarrow||\mathbb{K}_{1}-\mathbb{K}_{2}||_{2}$ (8)
@@ IV-B1 Dataset generation for network model
While collecting the dataset of image states $\mathbb{K}$ in subsection IV-A, we recorded the velocity (Eq. 5) applied to transition between consecutive frames. Using this data, we created a new dataset that includes pairs of consecutive image states ($\mathbb{K}_{1}$, $\mathbb{K}_{2}$) and their estimated joint displacements. This is calculated by multiplying the recorded velocity with the frame rate at which $\mathbb{K}$ was captured, as described in Eq. 6.
Please note here, only pairs of $\mathbb{K}$ captured in consecutive image frames are included in this data generation process. However, for constructing the graph $G$, it is essential to compute distances between arbitrary $\mathbb{K}$ pairs in $I$-space. To achieve this, a neural network is employed to predict the joint displacement across different pairs of image-states $\mathbb{K}$.
@@ IV-B2 Neural Network for Learned Distance Metric
To derive Eq. 7, we design a simple neural network using the aforementioned dataset to learn a distance more similar to the joint space distance. The model takes concatenated arrays of the starting image state ($\mathbb{K}_{\text{start}}$) and the subsequent image state ($\mathbb{K}_{\text{next}}$) as input and predicts the estimated joint displacement.
@@
To check for collisions in our model-free system, we employ an image-based polygon collision-checking framework. Polygons are formed by connecting pairs of consecutive keypoints (($\mathds{k}_{n-1}$), ($\mathds{k}_{n}$)) in one image state ($\mathbb{K}_{1}$) with their corresponding keypoints in the neighboring image state ($\mathbb{K}_{2}$). In Fig. 4, the polygons created by ($\mathds{k}_{1}$, $\mathds{k}_{2}$), ($\mathds{k}_{2}$, $\mathds{k}_{3}$), ($\mathds{k}_{3}$, $\mathds{k}_{4}$) and ($\mathds{k}_{4}$, $\mathds{k}_{5}$) of the start image state ($\mathbb{K}_{1}$) in green and its neighboring image state ($\mathbb{K}_{2}$) in red are examples of such polygons.
Each polygon is then checked for intersections with the obstacle boundaries defined by a safety margin which we consider to account for controller uncertainty. This polygon-based approach ensures a thorough and robust collision-checking process, effectively handling obstacles of any size. In Fig. 4 illustrates an example where the obstacle, enclosed by the red safety margin, intersects with the polygons defined by ([$\mathds{k}_{2}$, $\mathds{k}_{3}$],[$\mathds{k}_{3}$, $\mathds{k}_{4}$] and [$\mathds{k}_{4}$, $\mathds{k}_{5}$]), demonstrating the effectiveness of this method. The collision checking process and A* search is shown in A* Search with Lazy Collision Checking of Fig. 3.
Figure 4: Illustration of polygon-based collision checking for vision-only motion planning. In this image $4$ polygons (bordered with light blue) are defined by connecting consecutive keypoint pairs ($\mathds{k}_{1}$, $\mathds{k}_{2}$), ($\mathds{k}_{2}$, $\mathds{k}_{3}$), ($\mathds{k}_{3}$, $\mathds{k}_{4}$) and ($\mathds{k}_{4}$, $\mathds{k}_{5}$) of the image state ($\mathbb{K}_{1}$) in green and its neighboring image state ($\mathbb{K}_{2}$) in red. The yellow obstacle enclosed by the purple safety margin is situated right on the polygons formed by ([$\mathds{k}_{2}$, $\mathds{k}_{3}$],[$\mathds{k}_{3}$, $\mathds{k}_{4}$] and [$\mathds{k}_{4}$, $\mathds{k}_{5}$]), rendering these two configurations ineligible for a collision-free path.

### v2

@@ Abstract
This work presents a motion planning framework for robotic manipulators that computes collision-free paths directly in image space. The generated paths can then be tracked using vision-based control, eliminating the need for an explicit robot model or proprioceptive sensing. At the core of our approach is the construction of a roadmap entirely in image space. To achieve this, we explicitly define sampling, nearest-neighbor selection, and collision checking based on visual features rather than geometric models. We first collect a set of image space samples by moving the robot within its workspace, capturing keypoints along its body at different configurations. These samples serve as nodes in the roadmap, which we construct using either learned or predefined distance metrics. At runtime, the roadmap generates collision-free paths directly in image space, removing the need for a robot model or joint encoders. We validate our approach through an experimental study in which a robotic arm follows planned paths using an adaptive vision-based control scheme to avoid obstacles. The results show that paths generated with the learned-distance roadmap achieved 100% success in control convergence, whereas the predefined image space distance roadmap enabled faster transient responses but had a lower success rate in convergence.
@@ I Introduction
The goal of our research is to push the boundaries of purely vision-based control and motion planning for robotic manipulators, by decreasing reliance on explicit robot modeling or proprioceptive sensing. In doing so, we strive to use natural visual features along the robot’s body in image space, without attaching any external markers. We use multiple features along the robot’s body to uniquely distinguish full-body configurations that have the same end effector position, a characteristic common in redundant robots such as continuum manipulators. While the works in [15, 16, 17] provided algorithms to track natural keypoints and use them to achieve vision-based model-free control with decent transient responses, these algorithms are only designed to run in obstacle-free space and do not provide motion planning capability to avoid obstacles in the scene.
@@
In this work, we tackle the problem of collision-free path planning for purely vision-based robot control: we develop a planning and control scheme that does not rely on a robot model or on-board sensors (e.g. joint encoders) in runtime. As such, this strategy is especially useful for (but not limited to) soft/underactuated/inexpensive robots that are hard to model and may not have reliable proprioceptive sensing.
@@
This dataset was used to develop a simple neural network that approximates joint displacements based on keypoint locations in image space. These distance metrics were integrated into the roadmap construction. Once the roadmaps are constructed, polygon-based collision checking and A* [19] search are employed to ensure collision-free paths. These two approaches have different implications for vision-based robot control. In a nutshell, we observed that utilizing estimated joint distances in the roadmap results in smoother and more accurate tracking of the generated path, allowing the robot to stay in its defined workspace and avoid obstacles. In contrast, roadmaps based on Euclidean distances in image space can offer faster transient responses, albeit with potentially less reliable tracking. in environments with tightly spaced or irregularly shaped obstacles, these image space roadmaps frequently fail to generate feasible paths due to their poor alignment with the underlying joint-space feasibility. Our experiments demonstrate these aspects both in the presence and absence of obstacles.
@@ II Related Work
Among related works, the approaches proposed in [29] and [30] are the most closely aligned with ours. In [29], the authors plan trajectories by learning a low-dimensional latent space using three neural networks: an autoencoder, a latent dynamics model, and a collision checking network. This latent space allows planning without directly operating in the high-dimensional input space (e.g., raw images or complex joint configurations). While effective, their method requires extensive simulation-based training, which introduces a significant sim-to-real gap. Moreover, the system is validated only in simulation, and the reliance on multiple learned modules increases complexity, making real-world deployment more challenging. In contrast, our method operates directly in image space using visually tracked keypoints, avoiding the need for encoding, decoding, or learned dynamics. It uses real robot data exclusively, both for training and evaluation, and does not rely on simulation. This results in a lightweight, practical framework that is easier to implement and deploy in real-world scenarios. In [30], the authors learn the topology of a perceptual control manifold (PCM) using a topology-preserving neural network (TRN) constructed over both robot joint configurations and corresponding image features. This approach explicitly relies on access to joint space information and was designed specifically for a pneumatically driven robotic arm, limiting its generalizability. In contrast, our method requires only a dataset of image features and does not depend on any explicit robot model or configuration information. As a result, our approach is more general and can be easily modified to both rigid and soft robotic manipulators.
@@ III-A Model-based Motion Planning
$\footnotesize\mathcal{X}_{\text{obs}}=\{x\in\mathcal{X}\mid R(x)\bigcap\mathcal{O}\neq\emptyset\}$ (2)
where $R(x)\subseteq\mathbb{R}^{3}$ is the set of all points occupied by the robot at state $x$, and $\mathcal{O}\subset\mathcal{W}$ denotes the set of obstacles within the workspace [^1] This equation only encodes collisions; joint or kinematic constraints can be incorporated in a similar manner.. Let $\mathcal{X}_{\text{free}}=\mathcal{X}\setminus\mathcal{X}_{\text{obs}}$ denote the collision-free state space. Also, let $x_{\text{start}}\in\mathcal{X}_{\text{free}}$ and $\mathcal{X}_{G}\subseteq\mathcal{X}_{\text{free}}$ represent the start state and goal region, respectively. Motion Planning Problem: Given the motion planning tuple $(\mathcal{X},R,\mathcal{W},\mathcal{U},f,x_{\text{start}},\mathcal{X}_{G})$ find a time $\mathcal{T}$ and a set of controls $u:[0,\mathcal{T}]\rightarrow\mathcal{U}$ such that the motion described by Eq. 1 satisfies $x(0)=x_{\text{start}}$, $x(\mathcal{T})\in\mathcal{X}_{G}$ and $x(t)\in\mathcal{X}_{\text{free}}$.
@@
$\footnotesize I_{\text{obs}}=\{\mathbb{K}\in I|RI(\mathbb{K})\bigcap\mathcal{IO}\neq\emptyset\}$ (3)
@@
Here we note that even if the underlying function f is fully-integrable (holonomic-system) if $dim(I)>dim(\mathcal{X})$ the equivalent system dynamics equations $g$ will be non-holonomic as there will be paths in the higher dimensional image space $I$ that cannot be followed by the robot. In our approach we consider this issue, and propose a way to produce paths that approximately satisfy these constraints. In the first image of Fig. 2, the grey region illustrates the discretized representation of $I$, as each configuration is captured as an image frame, resulting in a finite set of $\mathbb{K}$ sampled at a fixed frame rate. The second image of Fig. 2 highlights similar discretized representation of $I_{\text{free}}$.
@@ IV Methodology
To compute the path, which is the main contribution of this work, we propose to use probabilistic road-map planner (prm) by adapting its subroutines to operate directly in $I$-space. Specifically, we opted to adapt the lazy-prm [33] planner due to its compatibility with our particular requirements. However, any sampling-based planner which relies on the same subroutines could be used.
@@
Similar to prm, lazy-prm operates in two phases: a building phase Alg. 1, where the roadmap is constructed without collision checking, and a query phase Alg. 2, where a path is searched and only the selected edges are checked for collisions. If any are invalid, the roadmap is updated and the search is repeated. This separation allows the roadmap to be built offline, before robot deployment, without considering obstacles, enabling reuse across different environments and physical setups.
We use the probabilistic roadmap method for its efficiency in multi-query scenarios, where a single precomputed graph supports multiple start–goal pairs. The lazy variant is particularly suitable as the environment remains mostly unchanged between planning queries. Its ability to handle high-dimensional spaces also makes it well-suited for our $I$-space representation based on visual keypoints. We first describe how lazy-prm works, followed by the modifications needed to adapt it to $I$-space.
The building-phase (Alg. 1) works with the following procedure. First, in line 4, a sample is generated and added in graph $G$. Then the k nearest neighbors are found by using a distance defined in $I$ (line 7) and are connected with edges. This continues until $N$ nodes are in the graph $G$.
@@
During the query-phase (Alg. 2) a new motion planning problem is solved. Given a $\mathbb{K}_{\text{start}}$ and $\mathbb{K}_{\text{goal}}$ they are added in the graph, and connected with their nearest-neighbors. Then a graph search algorithm e.g., A* is used to find a path. If edges of the path are in-collision they are updated accordingly in the roadmap, and the process repeats until a collision free-path is found. The three operations sample, k-nearest, coll_free, for Alg. 1 in lines 4, 7 and Alg. 2 typically require a model for the robot.
@@
sample: This function typically samples the configuration space uniformly. However, in the vision-only setting, we can’t directly sample in $I$-space as we don’t have the model of the robot. Since $dim(I)>dim(\mathcal{X})$ the keypoint vectors (image state $\mathbb{K}$) that correspond to actual configurations of the robot, will lie in a lower-dimensional (equal to $dim(\mathcal{X})$) manifold in $I$-space. Thus randomly sampling $I$-space will have 00 probability of sampling a valid configuration that lies on the valid manifold [23]. We describe how to address this issues in Section IV-A, by collecting and storing valid samples directly from the real robot.
k-nearest: This function usually relies on a distance defined in $\mathcal{X}$ and finds the nearest configurations that can be connected. However, since our representation in $I$-space is now a non-holonomic system, defining this distance is very challenging, as a straight line path in $I$-space defined by a simple Euclidean distance, might not be accurately followable by a controller. To mitigate this, we describe a learning-based approach to estimate the unknown joint-distance, in Section IV-B.
coll_free: This function checks if there is a collision for an edge in $\mathcal{X}$. Again this typically requires the model $R(x)$ for the robot. In the vision only case, we propose simple yet successful method to do collision checking with an $RI(x)$ directly in $I$-space Section IV-C.
In the next section we describe our proposed method for each of the aforementioned subroutines. Fig. 3 visually describes all the steps of the visual motion planning framework.
@@
$\footnotesize v_{j}=\min\left(\frac{M_{j}}{res\cdot dt},v_{\text{max}}\right),$ (5)
where, $M_{j}$ is the motion range or difference of limits of joint $j$, $res$, is the number of discrete steps used to traverse $M_{j}$, $dt$ is the duration allocated to complete each step, and $v_{\text{max}}$ is the maximum allowable velocity for joint $j$. By dividing the motion range of each joint into uniform increments based on $res$, we ensure that the robot systematically explores all possible configurations within its workspace. The resulting dataset of image states $\mathbb{K}$ is thus evenly distributed across the image space $I$, enabling robust coverage and accurate representation of the robot’s motion capabilities.
@@
$\footnotesize\mathbb{K}_{n}=K\cdot T_{cw}\cdot x_{n}\vspace{-0.75em}$ (6)
where $K$ is the camera intrinsic matrix, $T_{cw}$ is the camera extrinsics matrix derived from calibration processes described in [34, 35], and $x_{n}$ is the 3D robot configuration in workspace. This transformation projects the 3D configuration $x_{n}$, into their corresponding 2D (pixel) projection in the image described in [36], resulting in a set of $\mathds{k}$ for each image state $\mathbb{K}_{n}$ The process captures the robot’s motion across its visible workspace and creates a comprehensive dataset of $\mathbb{K}$, representing close to all feasible configurations in image space. Dataset of $\mathbb{K}$ can also be collected by following the process describe in [17].
The Keypoint Dataset Samples section of Fig. 3, represents this part of the workflow. The grey area in the left image of Fig. 2 illustrates how $\mathbb{K}$ are distributed within image space with each frame consisting of a set of $\mathds{k}$ or keypoints. This collection process yields a dataset of $\mathbb{K}$ that can be used to enable efficient roadmap construction for vision-based motion planning.
Since this data collection process relies only on observable image features, it can be easily applied to other robotic systems, including soft and rigid manipulators, enabling broad applicability of the proposed framework.
IV-B k-nearest: Learned and Image Space Metrics
@@
  • Learned distance, where the distance is learned by a neural network, trained to predict joint displacements between two image states. The input to the network is a pair of image states $\mathbb{K}_{1}$ and $\mathbb{K}_{2}$ and the output is an estimated joint displacement required to transition between them:
    $\footnotesize\text{dist}_{learned}()\leftarrow\text{NN}(\mathbb{K}_{1},\mathbb{K}_{2})$ (7)
@@
    $\footnotesize\text{dist}_{image}()\leftarrow||\mathbb{K}_{1}-\mathbb{K}_{2}||_{2}$ (8)
@@ IV-B1 Dataset generation for network model
While collecting the dataset of image states $\mathbb{K}$ in Section IV-A, we recorded the velocity (Eq. 5) applied to transition between consecutive frames. Using this data, we created a new dataset that includes pairs of consecutive image states ($\mathbb{K}_{1}$, $\mathbb{K}_{2}$) and their estimated joint displacements. This is calculated by multiplying the recorded velocity with the frame rate at which $\mathbb{K}$ was captured, as described in Eq. 6.
Only pairs of $\mathbb{K}$ captured in consecutive image frames are included in this data generation process. However, constructing the graph $G$, requires computing distances between arbitrary $\mathbb{K}$ pairs in $I$-space. To achieve this, a neural network is trained to predict joint displacements between different image-state pairs $\mathbb{K}$.
@@ IV-B2 Neural Network for Learned Distance Metric
To derive Eq. 7, we design a simple neural network using the aforementioned dataset to learn a distance more similar to the joint space distance. The model takes concatenated arrays of the starting image state ($\mathbb{K}_{\text{start}}$) and the subsequent image state ($\mathbb{K}_{\text{next}}$) as input and predicts the estimated joint displacement. We train the network using the Adam optimizer (learning rate: 0.005), a batch size of 32, and 400 epochs, minimizing mean squared error (MSE) loss. Performance is evaluated using RMSE, MAE, and $R^{2}$ metrics.
@@
To enable collision checking in our model-free system, we employ an image-based polygon collision-checking framework. Obstacles are defined directly in the image space by drawing contours around visible objects and expanding them with a safety margin to account for the robot’s volume and controller uncertainty. Each candidate path is checked by testing for intersection between the expanded obstacle region and polygons formed by corresponding keypoint pairs across two image states. This method effectively handles obstacles of arbitrary shape and size and does not require access to the robot’s model or workspace geometry The collision checking process and A* search is shown in A* Search with Lazy Collision Checking of Fig. 3.
Figure 4: Polygon-based collision checking in image space. Four polygons (light blue) are defined by connecting consecutive keypoints—($\mathds{k}_{1}$, $\mathds{k}_{2}$), ($\mathds{k}_{2}$, $\mathds{k}_{3}$), ($\mathds{k}_{3}$, $\mathds{k}_{4}$), and ($\mathds{k}_{4}$, $\mathds{k}_{5}$)—from the current image state ($\mathbb{K}_{1}$, green) to its neighbor ($\mathbb{K}_{2}$, red). A collision is detected if any polygon intersects the obstacle contour (yellow) inflated by a purple safety margin, as seen for ([$\mathds{k}_{2}$, $\mathds{k}_{3}$], [$\mathds{k}_{3}$, $\mathds{k}_{4}$], and [$\mathds{k}_{4}$, $\mathds{k}_{5}$]) in the first image and [$\mathds{k}_{4}$, $\mathds{k}_{5}$] in the shelf scenario.